OpenStudy (anonymous):

Find the length of the curve. \[y=1+6x^{3/2}\] 0=

OpenStudy (anonymous):

well the third line is how far I have gotten

OpenStudy (foolaroundmath):

try a substitution \(t^{2} = 1+81x\)

OpenStudy (anonymous):

or t=1+81x and dt=81 that actually solved it!!! thanks!

OpenStudy (foolaroundmath):

i personally try to avoid integrating square roots .. so i suggested t^2 .. t is also fine :)